UT 20140122 LUCI2 Nighttime Commissioning

Observer:JHeidt, APramskiy, DThompson
Telescope Operator: BGechetti
LBTO Support:

Plan for the night

Detector tests
Binocular presets

Summary

Lost 3/4 of night due to weather

Some MOS-alignment/spectroscopy

Detector-tests (linearity measurements)

Preparations

Details

All times are UT

Files are luci2.20140121.NNNN.fits
Afternoon darks
0001 0020 1x2.7399s LIR
0021 0040 1x6.8515s MER

Alexander corrected the filenaming to our UTdate rollover @noon

Files are luci2.20140122.NNNN.fits
Afternoon darks, continued
0006 0015 1x120.0s MER
0016 0026 1x300.0s MER
one extra 300s taken manually

W-unit Beamsplitter test for JC.
0027 @ -50 stage position (mm)
0028 @ -60 (SW limit)
0029 @ -40
0030 @ -4.4 (nomina centered pos'n)
0031 @ -60 (SW limit)

00:46 Sunset, weather: overcast, low wind, hum=5%, T=4deg

02:00 Field alignment, #32-36, dx,dy=-300,500

02:30 Alexander is testing some bino-presets track-mode

03:35 It cleared up a bit, its probably just a hole....lets git it a try, HH91, FeII #37, seeing oon images 0.5",
H2, #38, lost pointing/guiding, stopped, checked pointing, resend preset, failed, GS to faint given the cirrus we are facing

04:07 Ima movement., M67, AM=1.9, PA=0,90, #39/40, lost GS at PA=180, #41 crap, sent preset at PA=180 again, OK, PA=180/270 #42/43, PA=180 #44, clouds get thicker

05:06 Somewhat better now, S0952, GS selected from scritping tool not found, checked pointing, didn't help

05:32 HS1024, can't collimate, low AM and thick cirrus

05:48 S0952 again, using Alexanders command-line tool, works, sky #49, src #50, MOSview #51, ref.stars1-6 automatically found, but had problems to find all alignment stars, aligned, verification #52, OK, 150Ks@2.165 #53-63, 210H@1.62mue #64-73, lost GS at the 10th exp., #73 not useful since clouds are to heavy now.
Seeing on guider 0.5" all time

08:00 Telluric S0952, no way, clouds too thick now, we close for detector tests

08:23 Set up for Linearity data
0075 @TopRight corner ~1235 ADU = 62k in 140s
0076 0077 commit just first item...seems OK

08:32 Linearize_140s.xml (commit 100) All: J_high LIR
0078 0235 pairs of ramped exposures w/2.75s between
0236 0393 repeated full sequence

Calibrations

For ease of data reduction since detector is tilted

CalUnit? in, all images LIR, all normal
Spectro sieve, 10*10sec, G210@1.65mue, clear+H, halo2,#394-403
Spectro sieve, 1*10sec, G210@1.65mue, clear+H, lamp off, #404
Spectro sieve, 10*10sec, G210@1.25mue, clear+J, halo2,#405-414
Spectro sieve, 1*10sec, G210@1.25mue, clear+J, lamp off, #415
Spectro sieve, 10*10sec, G210@0.95mue, clear+z, halo2,#416-425
Spectro sieve, 1*10sec, G210@0.95mue, clear+z, lamp off, #426

S0952

Mask S0952M1, G150@2.165mue, clear+K, CalUnit? in, all images LIR, all normal
Lamp off, 5*2.75s, #427-431
Halo2 on, 5*10sec, #432-436
Lamps off, 5*10sec, #437-441
Argon on, 1*10sec, #441
Neon on, 1*10sec, #442
Xenon on, 1*10sec, #443
Lamps off, 5*10sec, #444-448

Mask S0952M1, G210@1.62mue, clear+H, CalUnit? in, all images LIR, all normal
Lamp off, 5*2.75s, #449-453
Halo2 on, 5*10sec, #454-458
Lamps off, 5*10sec, #459-463
Argon on, 1*10sec, #464
Neon on, 1*10sec, #465
Xenon on, 1*10sec, #466
Lamps off, 5*10sec, #467-471

End of Com#3 LUCI2

-- JochenHeidt - 22 Jan 2014

Topic revision: r5 - 22 Jan 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback